A heavy box is slid across a rough floor with an initial speed of 4 m/s. It stops moving after 8 sec. If the average resisting force of friction is 10N the mass of box in kg is

Grade:11

1 Answers

Ashish kumar sharma
6 years ago
According to the question-u=4m/sv=0m/st=8 secSo let`s first try to find out the accelerationHere`s how we can:Using v=u+atWe get, 0=4+a(8) a=1/2 m/s²So next step,Let frictional force to be FThus, F=m×a 10= m×1/2 So, m=20kgTherefore the mass is 20kg.Hope it helped:)